OpenStudy (anonymous):

Solve the quadratic equation. (x - 5)2 = 9

OpenStudy (anonymous):

is (x-5) raised to the 2nd?

OpenStudy (anonymous):

Take the square root of (x-5)^2 and the square root of 9. This results in x-5=3 Add 5 to both sides. x=8

OpenStudy (anonymous):

\(x-5=\sqrt{9}\\x=-3+5 \ or\ x=3+5\) remember that when calculating for the square root, it is either "positive" or "negative"